# OmniMER: Indonesian Multimodal Emotion Recognition via Auxiliary-Enhanced LLM Adaptation

### v1

@@ Abstract
Indonesian, spoken by over 200 million people, remains underserved in multimodal emotion recognition research despite its dominant presence on Southeast Asian social media platforms. We introduce IndoMER, the first multimodal emotion recognition benchmark for Indonesian, comprising 1,944 video segments from 203 speakers with temporally aligned text, audio, and visual annotations across seven emotion categories. The dataset exhibits realistic challenges including cross-modal inconsistency and long-tailed class distributions shaped by Indonesian cultural communication norms. To address these challenges, we propose OmniMER, a multimodal adaptation framework built upon Qwen2.5-Omni that enhances emotion recognition through three auxiliary modality-specific perception tasks: emotion keyword extraction for text, facial expression analysis for video, and prosody analysis for audio. These auxiliary tasks help the model identify emotion-relevant cues in each modality before fusion, reducing reliance on spurious correlations in low-resource settings. Experiments on IndoMER show that OmniMER achieves 0.582 Macro-F1 on sentiment classification and 0.454 on emotion recognition, outperforming the base model by 7.6 and 22.1 absolute points respectively. Cross-lingual evaluation on the Chinese CH-SIMS dataset further demonstrates the generalizability of the proposed framework. The dataset and code are publicly available. https://github.com/yanxm01/INDOMER
@@ I Introduction
Emotion recognition is central to affective computing and supports a wide range of applications, from mental health monitoring [1] to social media content moderation [2, 3]. As multimodal data becomes increasingly available, multimodal emotion recognition (MER), which integrates text, audio, and visual signals, has gained significant attention [4, 5]. Text conveys explicit semantic content, audio captures prosodic and vocal cues, and video encodes facial expressions and head movements. By combining these complementary modalities, MER systems capture the complexity of human affect more comprehensively than unimodal approaches [6, 7, 8]. However, the effectiveness of MER depends heavily on the availability of high-quality annotated data, which remains scarce for most languages beyond English.
Despite significant progress, existing MER research remains heavily English-centric. Benchmark datasets such as IEMOCAP [9], CMU-MOSEI [10], and MELD [11] have driven methodological advances, yet they predominantly feature English speakers in acted or scripted scenarios, limiting their capacity to generalize to spontaneous, cross-cultural emotional expressions. While CMU-MOSEAS [12] extends coverage to Spanish, Portuguese, German, and French, these remain relatively high-resource European languages. A critical gap persists for languages in Asia, Africa, and other underrepresented regions, where both annotated multimodal data and pretrained linguistic resources are scarce (Table I).
@@
Beyond data scarcity, MER in low-resource settings faces two compounding challenges. The first is cross-modal inconsistency. Emotional signals across modalities often conflict due to cultural communication norms [16, 17]. Indonesian speakers, influenced by values of social harmony (Rukun) and politeness (Sopan Santun), often use indirect or neutral language in text while displaying stronger emotional cues through vocal prosody or facial expressions [18]. Such modality conflicts make it difficult for models to learn consistent emotional representations. The second challenge is long-tailed class distributions. Natural emotional communication skews heavily toward neutral and positive expressions, while emotions like fear and disgust are rare [19]. This imbalance biases standard training procedures toward majority classes and amplifies spurious correlations, where models rely on dataset artifacts rather than genuine emotional signal [20]. With limited training data, these issues are further exacerbated, as models tend to overfit superficial patterns rather than learn robust multimodal representations [21, 22].
Recent advances in omni-modal large language models (LLMs) offer a promising direction for addressing these challenges. Models such as GPT-4o [23], Gemini [24], and Qwen2.5-Omni [25] unify text, audio, and visual understanding within a single architecture, demonstrating strong cross-modal perception capabilities. Their pretrained knowledge may help mitigate the limitations of small-scale datasets. However, directly applying omni-modal LLMs to emotion recognition without task-specific adaptation often yields suboptimal results, as the models may rely on superficial patterns rather than genuine emotional cues [26]. This motivates the need for targeted adaptation strategies that strengthen emotional grounding within each modality before multimodal fusion.
@@
  • We introduce IndoMER, the first multimodal emotion recognition benchmark for Indonesian, comprising 1,944 video segments from 203 speakers with temporally aligned text, audio, and visual annotations across seven emotion categories.
@@
  • Through extensive experiments on IndoMER, we demonstrate that OmniMER outperforms the base model by 7.6 and 22.1 absolute points in Macro-F1 on sentiment classification and emotion recognition respectively, with auxiliary tasks providing notable gains for minority emotion categories. Cross-lingual evaluation on CH-SIMS further validates the generalizability of the framework.
@@ II-B Low-Resource Multimodal Learning
To address data scarcity, recent work has explored transfer learning and cross-lingual adaptation by leveraging high-resource language models [52, 53, 54]. However, these approaches face fundamental limitations: morphological and syntactic mismatches across language families, culturally specific emotional expressions that fail to transfer, and the absence of basic NLP infrastructure for many low-resource languages [55]. These observations motivate the need for both dedicated datasets and specialized learning strategies tailored to low-resource multimodal emotion recognition.
@@ II-C Auxiliary-task Learning
Auxiliary-task learning provides additional supervisory signals that enhance representation learning [56], offering an effective strategy for adapting large pretrained models to specialized tasks. In multimodal emotion recognition, auxiliary tasks have been designed for individual modalities to capture emotion-relevant features. For the textual modality, word-level polarity annotations help models identify emotion-critical vocabulary [57]. For the visual modality, frame-level facial expression recognition guides the learning of emotion-aware visual representations [58]. For the acoustic modality, speech attributes such as pitch, speaking rate, and tone can be converted into natural language descriptions to enable LLM-based reasoning [59]. Additionally, hybrid strategies combining joint and multi-stage auxiliary learning have proven effective in conversational emotion recognition [60, 61].
Despite these advances, most existing auxiliary-task approaches are designed for high-resource settings and primarily focus on enhancing a single modality in isolation. In low-resource multimodal scenarios, where data scarcity coexists with cross-modal inconsistency, models are prone to learning spurious correlations from any single modality. A unified auxiliary learning strategy that jointly strengthens emotional representations across all modalities before fusion remains underexplored. This gap motivates our proposed OmniMER framework, which employs complementary auxiliary tasks across text, audio, and video to improve unimodal emotional grounding before multimodal fusion.
@@ III-A Data Collection
We collected individual monologue videos from publicly available social media platforms, such as YouTube and TikTok, prioritizing natural discourse where speakers exhibit emotional expressions through language, vocal prosody, and facial dynamics. To ensure topic diversity and mitigate potential biases from single-domain focus, we conducted systematic searches covering 13 themes: bloggers, books, celebrities, cooking, family, health, makeup, personal opinions, politics (non-controversial discussions), products, sharing, society, and tutorials (Fig. 1).
We applied the following selection criteria to maintain data quality and ethical standards: (1) each video contains only a single primary speaker to ensure clear modality alignment; (2) videos avoid sensitive topics related to religion, race, violence, or other controversial subjects; (3) content complies with platform community guidelines and contains no harmful, offensive, or politically inflammatory material; and (4) all videos are publicly available with no violation of privacy or intellectual property rights. Following these criteria, we obtained 207 source videos from 203 distinct speakers, ensuring both speaker diversity and data integrity. The gender distribution is approximately 40% male and 60% female, reflecting natural variation in social media content creators.
@@ III-B Segmentation and Annotation
Segmentation and annotation were conducted by seven Indonesian speakers and one Indonesian language expert to ensure linguistic accuracy and cultural relevance. Videos were segmented based on speaker pauses and content boundaries. For each segment, manual transcription was performed with careful preservation of grammatical features and word usage patterns of natural speech, rather than normalizing to formal written Indonesian. The language expert reviewed all transcriptions for accuracy and consistency, with special attention to regional linguistic variations.
Annotation proceeded in two phases. In the first phase, sentiment annotations were assigned to each segment on a scale from 1-1 (negative) to +1+1 (positive), with 0 representing neutral. Each segment was independently labeled by five Indonesian speakers, and the final sentiment label was determined via majority voting. In cases of disagreement, the language expert adjudicated based on cross-modal evidence, considering both linguistic content and non-verbal cues such as tone and facial expressions. In the second phase, emotion annotations were conducted using Ekman’s six basic emotions plus a neutral category, yielding seven categories: fear, disgust, anger, sadness, happiness, surprise, and neutral. Each segment received independent ratings from five Indonesian speakers and one language expert using a Likert scale from 0 to 3 (0 = no sign, 1 = weak, 2 = moderate, 3 = strong). The final emotion label was determined by the highest cumulative score across annotators. In ambiguous cases, the language expert resolved disagreements by considering contextual and emotional nuances.
@@ III-D Feature Extraction
To facilitate research using traditional multimodal fusion methods, we provide pre-extracted features for each modality following established practices in the MER literature. For the textual modality, we employ IndoBERT [^1] https://huggingface.co/indobenchmark/indobert-base-p1, a pre-trained language model specifically trained on Indonesian corpora. Token sequences are standardized to a fixed length of 69 tokens through padding or truncation, yielding 768-dimensional embeddings per token. For the acoustic modality, we utilize OpenSMILE [62] with the Geneva Minimalistic Acoustic Parameter Set (GeMAPS) [63] to extract 18-dimensional Low-Level Descriptors (LLDs) from audio sampled at 22,050 Hz. These features capture prosodic information including frequency, energy, and spectral characteristics. Acoustic sequences are standardized to 2,252 frames. For the visual modality, we employ OpenFace 2.2222https://github.com/TadasBaltrusaitis/OpenFace to extract 673-dimensional features from video frames at 25 fps, including facial action units, 2D/3D landmarks, head pose, and eye gaze. Visual sequences are standardized to 723 frames.
@@ IV OmniMER Framework
Figure 3 illustrates the proposed OmniMER framework. Built upon Qwen2.5-Omni, an omni-modal LLM that natively supports text, audio, and visual inputs, OmniMER introduces modality-specific auxiliary perception tasks to strengthen unimodal emotional grounding prior to multimodal fusion. The key insight is that directly fusing multimodal features without ensuring the quality of unimodal representations can propagate noise and spurious correlations, especially in low-resource settings. To address this, OmniMER employs three auxiliary tasks that explicitly supervise emotion-relevant cue extraction in each modality: emotion keyword extraction for text, facial expression analysis for visual, and prosody analysis for audio. While the main task uses both sentiment and emotion labels, the auxiliary tasks operate under weaker supervision using only modality-level sentiment annotations. The supervision signals are automatically constructed by prompting the omni-modal backbone to generate modality-specific emotional interpretations, which are then validated against sentiment labels. OmniMER consists of two core components: (1) modality-specific auxiliary perception tasks implemented via instruction-conditioned prompting, and (2) complementary training strategies, including multi-stage and hybrid optimization, to integrate auxiliary supervision with the main task.
@@ IV-B Auxiliary Tasks for Unimodal Emotional Grounding
OmniMER introduces auxiliary perception tasks for text, video, and audio modalities to strengthen unimodal emotional grounding before fusion. Unlike prior auxiliary-task approaches that typically focus on a single modality or require additional manual annotations [57, 58], our design has two distinguishing features. First, we employ a unified instruction-conditioned prompting paradigm across all three modalities, enabling consistent implementation within the omni-modal LLM framework. Second, the auxiliary supervision is automatically constructed using only existing modality-level sentiment labels, avoiding the need for fine-grained annotations such as word-level polarity or frame-level expressions.
Each auxiliary task is formulated as a modality-specific function that produces both a sentiment prediction and an interpretable explanation grounded in modality-specific evidence. The outputs are constrained to structured JSON format to ensure consistency and facilitate automatic validation.
@@ IV-B1 Text Auxiliary Task: Emotional Keywords Identification
Given a tokenized Indonesian utterance $T=\{w_{1},\dots,w_{n}\}$, the text auxiliary task aims to ground sentiment inference in emotion-related lexical cues by identifying salient keywords and phrases in the utterance:
@@
Prompt: Emotional Keywords Identification Instruction. You are an assistant for sentiment analysis of Indonesian text. Input. Text: {text} Task. Identify emotion-related keywords or phrases in the text and determine the overall sentiment supported by these cues. Return (JSON only).
@@ IV-C Auxiliary Supervision Construction
Although IndoMER provides sentiment annotations at both multimodal and modality-specific levels, it does not include fine-grained annotations describing modality-specific emotional evidence, such as emotion keywords in text, facial expressions in visual, or prosodic cues in audio. To obtain supervision for the auxiliary tasks, we leverage the generative capabilities of Qwen2.5-Omni with a two-step consistency-check mechanism (Fig. 4).
@@ V-C Sentiment- and Emotion-wise Performance Analysis
Overall, these class-wise results demonstrate that auxiliary tasks complement one another by strengthening unimodal emotional perception. The multi-stage strategy is particularly effective for recovering low-frequency emotions, whereas the hybrid strategy provides more stable improvements for majority categories. Despite these gains, recognizing extremely rare emotions such as disgust remains challenging, underscoring the difficulty of long-tailed recognition under limited supervision.

### v2

# Indonesian Multimodal Emotion Recognition via Auxiliary-Enhanced LLM Adaptation

@@ Abstract
Indonesian, spoken by over 200 million people, remains understudied in multimodal emotion recognition research despite its dominant presence on Southeast Asian social media platforms. To fill the gap, we introduce IndoMER, the first multimodal emotion recognition benchmark for Indonesian, which comprises 1,944 video segments from 203 speakers with temporally aligned text, audio, and visual annotations across seven emotion categories. The dataset poses realistic challenges for emotion recognition, including cross-modal inconsistency and long-tailed class distributions, which can be partly attributed to Indonesian culture. To address these challenges, we propose OmniMER, a multimodal adaptation framework built upon Qwen2.5-Omni that enhances emotion recognition through three auxiliary modality-specific perception tasks: emotion keyword extraction for text, facial expression analysis for video, and prosody analysis for audio. These auxiliary tasks help the model identify emotion-relevant cues in each modality before fusion, reducing the reliance on spurious correlations in low-resource settings. Experiments on IndoMER show that OmniMER achieves 0.582 Macro-F1 on sentiment classification and 0.454 on emotion recognition, outperforming the baseline model by 7.6 and 22.1 absolute points, respectively. Cross-lingual evaluation on the Chinese CH-SIMS dataset further demonstrates the generalizability of the proposed framework. The dataset and code are publicly available. (https://github.com/yanxm01/INDOMER)
@@ I Introduction
Emotion recognition is central to affective computing and supports a wide range of applications, from mental health monitoring [1] to social media content moderation [2, 3]. As multimodal data becomes increasingly available, multimodal emotion recognition (MER), which integrates text, audio, and visual signals, has gained significant attention [4, 5]. Text conveys explicit semantic content, audio captures prosodic and vocal cues, and video encodes facial expressions and head movements. By combining these complementary modalities, MER systems capture the complexity of human affect more comprehensively than unimodal approaches [6, 7, 8]. However, the effectiveness of MER depends heavily on large amounts of high-quality annotated data, which remains untrue for Indonesian.
Despite significant progress, existing MER research remains heavily English-centric. Benchmark datasets such as IEMOCAP [9], CMU-MOSEI [10], and MELD [11] have driven methodological advances, yet they predominantly feature English speakers in acted or scripted scenarios, limiting their capacity to generalize to spontaneous, cross-cultural emotional expressions. Although CMU-MOSEAS [12] extends coverage to Spanish, Portuguese, German, and French, these remain relatively high-resource European languages. A critical gap persists for languages in Asia, Africa, and other underrepresented regions, where both annotated multimodal data and pretrained linguistic resources are scarce (Table I).
@@
Beyond data scarcity, MER in low-resource settings faces two compounding challenges. The first is cross-modal inconsistency. Emotional signals across modalities often conflict due to cultural communication norms [16, 17]. Indonesian speakers, influenced by values of social harmony (Rukun) and politeness (Sopan Santun), often use indirect or neutral language in text while displaying stronger emotional cues through vocal prosody or facial expressions [18]. Such modality conflicts make it difficult for models to learn consistent emotional representations. The second challenge is long-tailed class distributions. Natural emotional communication skews heavily towards neutral and positive expressions, while emotions such as fear and disgust are rare [19]. This imbalance biases standard training procedures towards majority classes and amplifies spurious correlations, where models rely on dataset artifacts rather than genuine emotional signals [20]. With limited training data, these issues are further exacerbated, as models tend to overfit superficial patterns rather than learn robust multimodal representations [21, 22].
Recent advances in omni-modal large language models (LLMs) offer a promising direction for addressing these challenges. Models such as GPT-4o [23], Gemini [24], and Qwen2.5-Omni [25] unify text, audio, and visual understanding within a single architecture, demonstrating strong cross-modal perception capabilities. Their pretrained knowledge may help mitigate the limitations of small-scale datasets. However, directly applying omni-modal LLMs to emotion recognition without task-specific adaptation often yields suboptimal results, since theses models may rely on superficial patterns rather than genuine emotional cues [26]. This motivates the need for targeted adaptation strategies that strengthen emotional grounding within each modality before multimodal fusion.
@@
  • We introduce IndoMER, the first multimodal emotion recognition benchmark for Indonesian, which comprises 1,944 video segments from 203 speakers with temporally aligned text, audio and visual annotations across seven emotion categories.
@@
  • Through extensive experiments on IndoMER, we demonstrate that OmniMER outperforms the baseline model by 7.6 and 22.1 absolute points, respectively, on Macro-F1 for sentiment classification and emotion recognition, with auxiliary tasks providing notable gains for minority emotion categories. Cross-lingual evaluations on CH-SIMS further validate the generalizability of the framework.
@@ II-B Low-Resource Multimodal Learning
To address data scarcity, recent work has explored transfer learning and cross-lingual adaptation by leveraging high-resource language models [52, 53, 54]. However, these approaches face fundamental limitations: morphological and syntactic mismatches across language families, culturally specific emotional expressions cannot be transferred, and a unified NLP infrastructure for many low-resource languages still lacks [55]. These observations motivate the need for both dedicated datasets and specialized learning strategies tailored to low-resource multimodal emotion recognition.
@@ II-C Auxiliary-task Learning
Auxiliary-task learning provides additional supervisory signals that enhance representation learning [56], offering an effective strategy for adapting large pretrained models to specialized tasks. In multimodal emotion recognition, auxiliary tasks have been designed for individual modalities to capture emotion-relevant features. For the textual modality, word-level polarity annotations help models identify emotion-critical vocabulary [57]. For the visual modality, frame-level facial expression recognition guides the learning of emotion-aware visual representations [58]. For the acoustic modality, speech attributes such as pitch, speaking rate, and tone can be converted into natural language descriptions to enable LLM-based reasoning [59]. Additionally, hybrid strategies combining joint and multi-stage auxiliary learning have proven to be effective in conversational emotion recognition [60, 61].
Despite these advances, most existing auxiliary-task approaches are designed for high-resource settings and primarily focus on enhancing a single modality in isolation. In low-resource multimodal scenarios, where data scarcity coexists with cross-modal inconsistency, models are prone to learning spurious correlations from any single modality. A unified auxiliary learning strategy that jointly strengthens emotional representations across all modalities before fusion remains underexplored. This gap motivates our proposed OmniMER framework, which employs complementary auxiliary tasks across text, audio and video to improve unimodal emotional grounding before multimodal fusion.
@@ III-A Data Collection
We collected individual monologue videos from publicly available social media platforms, such as YouTube and TikTok, prioritizing natural discourse where speakers show emotional expressions through language, vocal prosody, and facial dynamics. To ensure topic diversity and mitigate potential biases from single-domain focus, we conducted systematic searches covering 13 themes: bloggers, books, celebrities, cooking, family, health, makeup, personal opinions, politics (non-controversial discussions), products, sharing, society, and tutorials (Fig. 1).
We applied the following selection criteria to maintain data quality and ethical standards: (1) each video contains only a single primary speaker to ensure clear modality alignment; (2) videos avoid sensitive topics related to religion, race, violence, or other controversial subjects; (3) content complies with platform community guidelines and contains no harmful, offensive, or politically inflammatory material; and (4) all videos are publicly available without violation of privacy or intellectual property rights. Following these criteria, we obtained 207 source videos from 203 different speakers, ensuring both speaker diversity and data integrity. The gender distribution is approximately 40% male and 60% female, reflecting natural variations in creators social media contents.
@@ III-B Segmentation and Annotation
Segmentation and annotation were conducted by seven Indonesian speakers and one Indonesian language expert to ensure linguistic precision and cultural relevance. Videos were segmented based on speaker pauses and content boundaries. For each segment, manual transcription was performed with careful preservation of grammatical features and word usage patterns of natural speech, rather than normalizing to formal written Indonesian. The language expert reviewed all transcriptions for accuracy and consistency, with special attention to regional linguistic variations.
The annotation process was implemented in two phases. In the first phase, sentiment annotations were assigned to each segment on a scale from 1-1 (negative) to +1+1 (positive), with 0 representing neutral. Each segment was independently labeled by five Indonesian speakers and the final sentiment label was determined via majority voting. In cases of disagreement, the language expert adjudicated based on cross-modal evidence, considering both linguistic content and non-verbal cues such as tone and facial expressions. In the second phase, emotion annotations were conducted using Ekman’s six basic emotions plus a neutral category, yielding seven categories: fear, disgust, anger, sadness, happiness, surprise, and neutral. Each segment received independent ratings from five Indonesian speakers and one language expert using a Likert scale from 0 to 3 (0 = no sign, 1 = weak, 2 = moderate, 3 = strong). The final emotion label was determined by the highest cumulative score across annotators. In ambiguous cases, the language expert resolved disagreements by considering contextual and emotional nuances.
@@ III-D Feature Extraction
To facilitate research using traditional multimodal fusion methods, we provide pre-extracted features for each modality following established practices in the MER literature. For the textual modality, we employ IndoBERT [^1] https://huggingface.co/indobenchmark/indobert-base-p1, a pre-trained language model specifically trained on Indonesian corpora. Token sequences are standardized to a fixed length of 69 tokens through padding or truncation, yielding 768-dimensional embeddings per token. For the acoustic modality, we utilize OpenSMILE [62] with the Geneva Minimalistic Acoustic Parameter Set (GeMAPS) [63] to extract 18-dimensional Low-Level Descriptors (LLDs) from audio sampled at 22,050 Hz. These features capture prosodic information including frequency, energy and spectral characteristics. Acoustic sequences are standardized to 2,252 frames. For the visual modality, we employ OpenFace 2.2222https://github.com/TadasBaltrusaitis/OpenFace to extract 673-dimensional features from video frames at 25 fps, including facial action units, 2D/3D landmarks, head pose, and eye gaze. Visual sequences are standardized to 723 frames.
@@ IV OmniMER Framework
Figure 3 illustrates the proposed OmniMER framework. Built upon Qwen2.5-Omni, an omni-modal LLM that supports text, audio, and visual inputs, OmniMER introduces modality-specific auxiliary perception tasks to strengthen unimodal emotional grounding prior to multimodal fusion. The key insight is that directly fusing multimodal features without ensuring the quality of unimodal representations can propagate noise and spurious correlations, especially in low-resource settings. To address this, OmniMER employs three auxiliary tasks that explicitly supervise emotion-relevant cue extraction in each modality: emotion keyword extraction for text, facial expression analysis for visual, and prosody analysis for audio. While the main task uses both sentiment and emotion labels, the auxiliary tasks operate under weaker supervision using only modality-level sentiment annotations. The supervision signals are automatically constructed by prompting the omni-modal backbone to generate modality-specific emotional interpretations, which are then validated against sentiment labels. OmniMER consists of two core components: (1) modality-specific auxiliary perception tasks implemented via instruction-conditioned prompting, and (2) complementary training strategies, including multi-stage and hybrid optimization, to integrate auxiliary supervision with the main task.
@@ IV-B Auxiliary Tasks for Unimodal Emotional Grounding
OmniMER introduces auxiliary perception tasks for text, video, and audio modalities to strengthen unimodal emotional grounding before fusion. Unlike previous auxiliary-task approaches that typically focus on a single modality or require additional manual annotations [57, 58], our design has two distinguishing features. First, we employ a unified instruction-conditioned prompting paradigm across all three modalities, enabling consistent implementation within the omni-modal LLM framework. Second, the auxiliary supervision is automatically constructed using only existing modality-level sentiment labels, avoiding the need for fine-grained annotations such as word-level polarity or frame-level expressions.
Each auxiliary task is formulated as a modality-specific function that produces both a sentiment prediction and an interpretable explanation grounded in modality-specific evidence. The outputs are constrained to the structured JSON format to ensure consistency and facilitate automatic validation.
@@ IV-B1 Text Auxiliary Task: Emotional Keywords Identification
Given a tokenized Indonesian utterance $T=\{w_{1},\dots,w_{n}\}$, the text that auxiliary task aims to ground sentiment inference in emotion-related lexical cues by identifying salient keywords and phrases in the utterance:
@@
Prompt: Identification of Emotional Keywords Instruction. You are an assistant for sentiment analysis of Indonesian text. Input. Text: {text} Task. Identify emotion-related keywords or phrases in the text and determine the overall sentiment supported by these cues. Return (JSON only).
@@ IV-C Auxiliary Supervision Construction
Although IndoMER provides sentiment annotations at both multimodal and modality-specific levels, it does not include fine-grained annotations describing modality-specific emotional evidence, such as emotion keywords in text, facial expressions in visual or prosodic cues in audio. To obtain supervision for the auxiliary tasks, we leverage the generative capabilities of Qwen2.5-Omni with a two-step consistency-check mechanism (Fig. 4).
@@ V-C Sentiment- and Emotion-wise Performance Analysis
Overall, these class-wise results demonstrate that auxiliary tasks complement each other by strengthening unimodal emotional perception. The multi-stage strategy is particularly effective for recovering low-frequency emotions, whereas the hybrid strategy provides more stable improvements for majority categories. Despite these gains, recognizing extremely rare emotions such as disgust remains challenging, underscoring the difficulty of long-tailed recognition under limited supervision.